C Band Gregorian Dome Receiver Log

as of Friday, 03-Dec-2021 07:07:01 AST

Go to Most Recent Entry

ANS anolla@naic.edu 10/7/99 Problem: There is a helium leak in the cryo system. Correction: A leaky hose was founded in the supply line at the rotary floor, hose was tightening up and quit leaking.
ANS anolla@naic.edu 1/11/2000 There was water condensed at the window of the recvr. Dewar .We recommend to add pressures to the horn to avoid this.
ANS anolla@naic.edu 2/4/2000 System shutted off due to contamination.Need crosshead replacement.
ANS anolla@naic.edu 2/7/2000 System cooled down.Crosshead ED 015 replaced with crosshead ED 013 due to contamination on cryo system.
ANS anolla@naic.edu 3/29/2000 New auto shutter installed. Pneumatic type developed by Jose Vives.
ANS anolla@naic.edu 4/17/2000 Pneumatic shutter removed due to problems with it. It was getting stocked due to faulty air compressor.
ANS anolla@naic.edu 6/8/2000 Auto shutter installed again , isolation was checked but is it too poor , just 20 db.
ANS anolla@naic.edu 9/28/2000 Receiver is hot due to a short ckt. At the cryo compressor (ED 001) The AC wiring at the power sw made a short ckt that melted a he line.
ANS anolla@naic.edu 10/10/2000 Receiver had a bad RF amp on chann. A. It was replaced, Amp sn on chann A is C11 & in chann B is C12.The sn of the bad amp on chnn A was C10.
ANS anolla@naic.edu 10/11/2000 Receiver cooled down.Chann A T Sys is 38 k & chann B is 32 k .
ANS anolla@naic.edu 10/25/2000 Receiver crosshead replaced due to contamination on it , it came from la large helium loss in comp.Crosshead ED 013 replaced with ED 015. Then it is been cool with comp 010 , so it will be three 350 crossheads with this comp.
ANS anolla@naic.edu 11/2/2000 After a week been run by comp. ED 010 the cryo temps are 12.4/61.7 k
ANS anolla@naic.edu 12/11/2000 Receiver cal box taken down for a check out . Carl Hailes complain that there is a 45 deg phase shift between Cal on channels, so Lisa is checking it ,
ANS anolla@naic.edu 12/13/2000 Receiver cal box was reinstalled , system has to be recalibrated.
ANS anolla@naic.edu 12/19/2000 Polarization check performed by WG & EC .Chan left is left and right is right.
ANS anolla@naic.edu 12/27/2000 CB cryo hooked to comp 008 .Was on comp 010.
ANS anolla@naic.edu 12/29/2000 The cryo was wrongly tie to comp 012 , it was suppose to go to comp 008 , the same comp used by 430 mhz.Was hooked back to comp 010.
ANS anolla@naic.edu 1/18/2001 The cryo comp 010 was replaced with 012, now is running with the 430 MHz Dewar.It was running with the NLB & NSB.
ANS anolla@naic.edu 2/19/2001 The cryo compressor was changed, it was 012 and was removed to 009, and it is in parallel with the 430.
ANS anolla@naic.edu 7/9/2001 The post amp chassis was taken down in order to check the noise diodes.
ANS anolla@naic.edu 7/10/2001 Post amp chassis box was installed back because it was going to be used tonight by Tapasi Goshi.No measurements were taken on it.
ANS anolla@naic.edu 7/16/2001 Post amp chassis box was taken down againg in order to check its cal noise diodes on it.
ANS anolla@naic.edu 7/30/2001 Cryo system turned off in order to replace contaminated crosshead.
ANS anolla@naic.edu 8/1/2001 Replaced crosshead on Dewar , it was ED 015 , now is ED 013. The 16 stage takes a much longer time to cool down due to the mass on its stage (ant construction) , you can have a temp of 45 k on the 70 stage after 4 hours , but at 16 can be 200.WATCH OUT FOR THESE .Also the supply cryo line (MSF 005) was replaced with a serviced one ( MSF 004).
ANS anolla@naic.edu 8/3/2001 C-band post amp modifications: ------------------------------ 1) The channel B cal noise diode output was found to vary with the TTL signal level and Channel A diode's initial turn ON time was found to be about 6 to 8 seconds. Pending repairs at Micronetics factory, an interface circuit with a voltage regulator (LM 317LZ) has been built and incorporated. Now, the TTL signal turns the 5V-regulated output ON and OFF, ensuring the stability of diode B noise output. The turn-on time problem with diode A remains. Both these diodes will be sent for repair after the VSOP runs in August. 2) The heat sinking of the post amp was insufficient. It was improved by adding Al blocks with fins and properly fixing with heat sink compound.
ANS anolla@naic.edu 8/8/2001 The ED 009 cryo comp was taken down for service, so ED 011 was hooked to the system.
ANS anolla@naic.edu 8/13/2001 The ED 011 cryo comp removed from the system in order to hook up it's proper one to it, these is comp ED 012. Also we suspect that comp ED 011 is leaking.
ANS anolla@naic.edu 11/9/2001 The heliax antennas to measure polarization cover from 4.5 to 5.5 Ghz , so cf is 5 Ghz.(L Wray measured them).
ANS anolla@naic.edu 11/26/2001 System was warmed due to a power dip from yesterday. JR & IC went up to restarted.
ANS anolla@naic.edu 1/9/2002 System was warmed up in order to replace its crosshead due to contamination, it was cycling. Crosshead was ED 013 , now is ED 019.
ANS anolla@naic.edu 1/24/2002 There is an slope in the bandpass , LW working on it.
ANS anolla@naic.edu 4/17/2002 System at room temperature , cryocomp quit (012).
ANS anolla@naic.edu 4/24/2002 System was cooled down , cryocomp is now ED 007. Also there is a problem with the Tsys on chann A ,it has a Tsys of about 50 k instead 32 k , chann B is ok.
ansanolla@naic.eduTuesday, August 13, 2002 15:13:48Post amp box down during the dual beam run to work on it.
LSWlwray@naic.eduFriday, August 16, 2002 16:09:10Post amp box still down, the receiver is not useable without it.
LSWlwray@naic.eduFriday, August 23, 2002 16:41:27Post amp chassis was put back on the telescope today. There were some wiring problems with the noise diode regulation circuit, but these have been fixed. The receiver is back in working condition.
ansanolla@naic.eduWednesday, August 28, 2002 09:49:26Focal point of the receiver readjusted , new shims on Dewar. They are 1.08".PP checking its performance at the new focal point.
ansanolla@naic.eduMonday, September 9, 2002 11:52:09 The CB shutter was not workink due to IC HSB shutter driver had something wrong with it and both TTL shutters driving signals run in parallel.HSB removes and everything worked ok.
ansanolla@naic.eduWednesday, September 25, 2002 11:48:08Cryocomp down due to power dip. Dewar warmed up , been cooled.
lswlwray@naic.eduWednesday, October 16, 2002 11:40:49Found HEMT LED switch for Ch A was off. This increased the Tsys from 33 to 50K. It has been switched back on.
lswlwray@naic.eduSaturday, November 2, 2002 11:40:22The kapton cover over the horn was removed on Wednesday October 30. In its place some Flouroglas fabric (permeable to radio waves, not to water) was put on, for testing.
LSWlwray@naic.eduFriday, November 22, 2002 17:43:25Nov 22: added a SMA female-SMAfemale and a SMA male-SMAmale connector to the output of Ch B's postamp, to put in a small phase delay as needed by the VLBI people.
ans anolla@naic.eduWednesday, February 26, 2003 11:25:45CB post amp chassis down in order to replace the cal diodes.
lswlwray@naic.eduMonday, March 3, 2003 16:28:37installed post amp box for pointing tests. old noise diodes still in. Changed attenuators after post amp from: 0/0 dB (A/B) to 3/2 dB, rebalancing levels after upstairs RF switch change.
lswlwray@naic.eduMonday, April 14, 2003 16:35:08removed post amp chassis in order to install new diodes.
ansanolla@naic.eduTuesday, April 29, 2003 08:43:16Chassis with new cal diodes back on 4/24/2003 . System was calibrated on 4/25/2003.
ansanolla@naic.eduTuesday, May 6, 2003 13:01:16There was a ripple on chann B due to a loose sma connector in the cal diodes area. Ganesh and Lisa found it.
ansanolla@naic.eduThursday, May 8, 2003 11:42:04For some reason the calibration was not working , we suspect of an intermittant TTL drive cable. Because it started working after we made several test.
ansanolla@naic.eduFriday, May 16, 2003 13:22:02P.P. calibrated the system this morning.
ansanolla@naic.eduWednesday, January 7, 2004 15:05:19The cryocomp ED 007 was replaced by the ED011 due to an oil leak on it.
ansanolla@naic.eduWednesday, January 14, 2004 13:07:11Cryocomp swap again due to a failure of cryocomp ED011.It was replaced by an ALFA one on 1/8/2004. But the they after it was found that the ALFA cryocomp was contaminated making both attached to it getting stock (430 & CB ). Then CB cryopump was purged an attached again to the Alfa cryocomp.
ansanolla@naic.eduTuesday, January 20, 2004 16:31:17The cryopump ED019 was replaced by ED015 . It was contaminated by the ALFA cryocomp.
ansanolla@naic.eduWednesday, January 21, 2004 16:14:56System back to operation. Was cooled down.
ansanolla@naic.eduTuesday, February 3, 2004 14:35:42System cryo temp is unstable , for some reason its performance quit.The crosshead was recently swap.
ansanolla@naic.eduTuesday, March 30, 2004 16:15:00There it was a short circuit at the rotary floor. For some reason also the main UPS (1400 VA) shorted out due to a shor at the remote sw for the meters inside cabinet #1. The remote sw was by passed.
ansanolla@naic.eduMonday, April 12, 2004 14:59:30There is not shuter due to burned out air compressor.
ansanolla@naic.eduTuesday, April 13, 2004 15:51:46Yesterday we found that the air compressor for the shutter was burned out. So today we replaced it with a new one. Shutters are working properly.
lswlwray@naic.eduThursday, June 10, 2004 16:27:40Monday May 31: Power failure in dome for four hours caused C Band to warm up. When it cooled back down, Ch A's Tsys read 12K higher than before (used to be 32K/32K, now 44K/32K).
lswlwray@naic.eduThursday, June 10, 2004 16:28:34Troubleshooting Tsys problem. Possibly a bad connector at input of ChA postamp box. Now Tsys is 38K/32K. Will bring down postamp box next week (dual beam) to investigate.
lswlwray@naic.eduTuesday, June 15, 2004 08:22:58June 14: Took down the postamp, biasbox, RF cables to dewar and cal cables to dewar, and interconnect cable between bias and postamp box. Will look for high Tsys problem.
lswlwray@naic.eduMonday, June 21, 2004 08:38:08June 18: Replaced postamp and bias chassis. The TTL cal cable from the IFLO cabinet to the postamp chassis was bad (weak connection at the postamp chassis end). Replaced with a semi-flex SMA cable.
lswlwray@naic.eduTuesday, June 22, 2004 13:56:45Troubleshooting downstairs: determined that the system sensitivity in Ch A is not higher than in Ch B, and that the cal value is not as high as we expect. For all cal types this is true. Therefore, problem points to losses in cal cable going to dewar.
lswlwray@naic.eduWednesday, June 23, 2004 16:25:54Performed many tests to figure out the Tsys problem (A 15K higher than B). 1. exchanged cal cables with lab cables. 2. put back original cal cables, then exchanged the RF cables between the dewar and postamp with lab cables. 3. put back original RF cables, then swapped the RF at the input of the postamp and also the output of the postamp. None of these tests showed any change in the tsys.
ansanolla@naic.eduFriday, July 2, 2004 11:33:35System down due to power shut down (main power line shorted at the rotary joint).
ansanolla@naic.eduSaturday, July 3, 2004 16:31:02System been cooled , will be ready for tonight.
ansanolla@naic.eduWednesday, August 25, 2004 11:45:22The 430 & CB cryocopmp replaced. It was ED 007 ,now is ED 012. The ED 007 down for prev. maint. & modifications.
ansanolla@naic.eduWednesday, September 22, 2004 14:40:22Cal cables were replaced , channel A was broken at the cal diode out. They are 21" long , also are phase match (6 deg at 5 ghz).
W Genterwgenter@naic.eduFriday, October 1, 2004 11:40:56C-Band receiver will be out of service from the evening of October 1 to October 8, 2004. This service is to repair a shorted amplifier LED.
ansanolla@naic.eduMonday, October 11, 2004 12:36:46System been cooled down.Exiter diode short fixed by BG.
ansanolla@naic.eduTuesday, October 12, 2004 13:12:22System cycling in the 16 deg stage. Suspect of the 16 deg seal.
W. Genterwgenter@naic.eduSunday, October 31, 2004 10:44:36This receiver is scheduled for a refrigerator change the week of 11/1/04. The refrigerator was last changed on 10/11/04. The Dewar temperature has remained at 20+k since cooled down on October 12, 2004.
ansanolla@naic.eduMonday, November 1, 2004 11:37:34System warmed up in order to replace crosshead.
ansanolla@naic.eduMonday, November 8, 2004 13:00:15The cryohead was replaced on 11/1/2004 , it was cooled down . Now 16stage is ok.Today the polarization box was added to the system , but still need the control bit to swap between linear and circular.Set at linear.
ansanolla@naic.eduTuesday, November 9, 2004 10:58:24System down in order to check the post & pol box phase shift.
ansanolla@naic.eduMonday, November 15, 2004 13:04:23System back to normal.Tsys is chan A 33.2 k and B 34.7 k.It was at lab for phase measurments.
ansanolla@naic.eduTuesday, November 30, 2004 14:30:28The calibration system was not working properly in all the receivers at the Dome . It was due that the post amp chassis for the CBH and XB were off. These was loading the other recvrs calibrators.
ansanolla@naic.eduTuesday, November 30, 2004 14:35:27The recvr channels were check for isolation with the hex antennas at the bowl there were a 17 db difference between channels. Was at the Fo of 4.95 Ghz.
ansanolla@naic.eduWednesday, January 19, 2005 13:19:14The noise injection board was replaced with the modified one to avoid oscillations. The one at the SBL was causing an rfi at 60 Mhz.
lswlwray@naic.eduThursday, January 27, 2005 12:45:54Jan 20: brought up new Polarization box but did not put into system yet. Jan 27: testing box with receiver. The box is not in the system yet. Need to make a control cable. CB is native linear, this box allows for circular pol.
ansanolla@naic.eduMonday, March 21, 2005 15:22:49Cryocomp replaced , it had a knocking sound.It was ED 012 , now is ED 007
ansanolla@naic.eduMonday, April 18, 2005 15:29:57Receiver down for modifications since last week.
ansanolla@naic.eduWednesday, May 25, 2005 13:12:51System back to service after replacing the rf amps at the Dewar with the indium phosphite ones (2 stages). Been cooled.
ansanolla@naic.eduThursday, May 26, 2005 11:04:01System cooled & operational.
ansanolla@naic.eduThursday, June 30, 2005 13:10:05The polarization box was installed in the receiver. Phil rechecked its command function and is ok.
W. Genterwgenter@naic.eduFriday, July 22, 2005 15:34:08Polorization check for VLBI run.
ansanolla@naic.eduFriday, August 26, 2005 13:06:58Due to power failure on 8/25/2005 system warmed up.Been cooled.
ansanolla@naic.eduTuesday, August 30, 2005 15:59:01System back to normal.
ansanolla@naic.eduWednesday, August 31, 2005 16:08:09Shutter was not working due that air compressor was bad. Replace it.
ansanolla@naic.eduTuesday, September 13, 2005 10:12:23Shutter plate cover fall out after the dual beam run.Need locktite on the holding screw.
ansanolla@naic.eduFriday, October 28, 2005 14:16:10The system was warmed out in order to replace its crosshead.
ansanolla@naic.eduMonday, October 31, 2005 14:54:02Crosshead replaced it was ED 005,now is ED 014.
ansanolla@naic.eduTuesday, November 1, 2005 09:55:09System cool , 16 stage is reading 14k.
W. Genterwgenter@naic.eduWednesday, December 14, 2005 08:35:19Reported drop-out in both channels during observations. Checked 10GHz converter LO and Power Supplies. All cables stable. Connectors tight.
ansanolla@naic.eduFriday, January 27, 2006 14:58:47System was calibrated for Chris & Tapasi.
ansanolla@naic.eduTuesday, May 9, 2006 14:35:31Cryo comp replaced , it was ED 007 , now is ED 015.
ansanolla@naic.eduTuesday, May 16, 2006 11:35:38CBL pol box removed for measurments until further notice.
ansanolla@naic.eduMonday, June 12, 2006 14:33:15Polarization check perfomed today (Carlitos in charge). Polarization box still downstairs
ansanolla@naic.eduMonday, June 12, 2006 14:36:45Forgot to mention that the polarization check out results were the same as before.
ansanolla@naic.eduThursday, July 6, 2006 12:51:09System calibrated today by Felipe.
ansanolla@naic.eduThursday, July 6, 2006 12:52:02System was calibrated today by Felipe.
ansanolla@naic.eduTuesday, January 23, 2007 11:48:08Polarization sw on it.Polarization can be adjusted from outside the chassis now.
ansanolla@naic.eduMonday, January 29, 2007 11:58:06Ganesh found that the gain in channel A is 3 db higher than B. Is affecting the pol test being performed on system.
ansanolla@naic.eduMonday, March 19, 2007 11:52:39System warmed up on 3/17/2007.BG turn off the crosshead that day.
ansanolla@naic.eduWednesday, March 21, 2007 16:24:56After replacing crosshead , system being cooled. Cryohead was ED 014 , now is ED 025.
ansanolla@naic.eduThursday, March 22, 2007 11:15:12System back to normal.Dewar valve closed but vacuum pump still on , being cleaned.
ansanolla@naic.eduTuesday, April 24, 2007 10:41:36System down due to painting project.
CRVThursday, January 3, 2008 09:41:08Vacuum pump connected to the dewar and pumping on Dic.11th,2007
CRVThursday, January 3, 2008 09:42:14Dewar cooldown Started on Dic.11th,2007 afternoon.
CRVThursday, January 3, 2008 09:43:03Dewar was cold by Dic.12th,2007
crvTuesday, January 29, 2008 16:53:34The system warmed down during this weekend 430mhz dual beam run. The compressor braeker opened during thge power dips.
crvTuesday, January 29, 2008 16:56:53The system was vac. and cooled 1/29/08
crvMonday, August 10, 2009 10:33:47The dewar was warmed up due to the power failures of sunday august 2nd and wednesday august 5th 2009. The Dewar was vac. and started coolin on friday agust the 7th 2009.
crvcriosMonday, March 22, 2010 14:26:10sytem temp. is raising aparantly due to displacer seal ploblems.
crvcrios@naic.eduThursday, March 25, 2010 16:26:15The cryo pump was replacedand started cooldown on march 24th 2010. today the vacc. pump was removed and the temp. was 16K=8.2K .
criosFriday, April 30, 2010 17:20:56System was warmed during the last week because of racheting sound. Reciever started cooling today on the new comp. C3 that was raised yesterday to replace the alfa compressor C2 that failed.
crvWednesday, June 2, 2010 10:34:59System was moved to comp ed013 after replacing the compressor absorver on May 28th,2010.
crvcrios@naic.eduMonday, October 25, 2010 11:28:01The shutter was reported to not close completely. The valve used to close the shutter was adjusted. The shutter was tested and closing fine.
crvcrios@naic.eduWednesday, December 8, 2010 13:53:53Monday, October 25, 2010 11:28:01 while checking the shutter we notice that the moister seal was missing we installed a new seal to the horn.
crvcrios@naic.eduMonday, May 14, 2012 10:43:57Dewar brought down today to repair Ch B. Oscillations.
Denisdurbain@naic.eduMonday, August 27, 2012 11:45:59The S-band high bias box was modified by Carlos to operate the C-band amplifiers. The Bias boxes were swapped to investigate the instability problem with the Channel-A stage-2 current.Phil is monitoring the bias and RF outputs.
Denis Urbaindurbain@naic.eduWednesday, August 29, 2012 10:09:59The bias cable was swapped. The shutter was closed for the duration of the maintenance time to eliminate possible RFI problem. The bias values are as always being monitored. The instability in the Canal A stage 2 remains after bias box swap 2 days ago.
criosWednesday, September 19, 2012 15:25:26The dewar was turned off and lowered to the shop for repair. Amp. channel A was replaced because of stability problems . Chalmers 4-8 GHz LNA SN(6h-3 18)replaced by SN(6h-3 19).
Denis Urbaindurbain@naic.eduThursday, October 11, 2012 08:12:17The pneumatic shutter was serviced, new tube and stainless steel fittings were installed.
dana whitlowdwhitlow@naic.eduThursday, October 18, 2012 17:12:11Yesterday we investigated the cause of the post-reinstallation gain and Tsys loss in Pol A with the receiver opened up in the dome, and found that the problem was apparently due to a bad SMA elbow and/or barrel in the path between the isolator and LNA. We replaced both suspect items, then proceeded to make room temp noise measurements for each of amplifiers 18 & 19 in Pol A. 19 looked better by several degrees, so we installed it. Note that this was a risk because 19 was the amplifier that exhibited the gain and bias instabilities which precipitated this whole sordid affair. But based on extensive logging of bias values without seeing any problem, and on the fact that we have learned something about the cause and cure for the problem (intermittent oscillations at ~955 MHz which can be cured by a minor tweak of Id), we decided to take the gamble. The receiver began cooling today (18 Oct) about 10AM, and is now down to ~75K with tsys values looking quite promising. The final word on stability is yet unknown.
Denis Urbaindurbain@naic.eduThursday, October 18, 2012 17:13:39The Channel A amplifier had been replaced a few weeks ago with the spare Chalmers amplifier #18. Since the swap the Tsys on Channel A has been about 15deg higher. The receiver was warmed-up and it was opened in the dome to investigate the cause of this higher Tsys. Gain measurements on both channels pointed out to a SMA elbow between the isolator and the amplifier. The elbow was replaced. Noise temperature done by Dana did show however a higher temperature on Channel A even after the new elbow was installed. The old amplifier #19 was installed and the temperature dropped to the same level as Channel B. We decided to leave the amplifier #19 in place and close the receiver. We will deal with the bias stability issue later if it shows up again.
Denis Urbaindurbain@naic.eduWednesday, November 7, 2012 13:11:55Pol A amplifier bias has some stability problem, unfortunately the problem is random or at least not constant. The pol B bias was adjusted to eliminate some out of the band oscillation. The some variation of the bias values did not affect the Receiver temperature. I adjusted the Pol A bias to match the Pol B. Vd1= 0.8V, Id1= 480mA, Vd2= 0.8V, Id2= 480mA.
Dana Whitlowdwhitlow@naic.eduMonday, February 4, 2013 15:15:32This morning Denis & I investigated the complaint of fairly large variations in noise level and of PolB stage 2 Id, both apparently synced to crosshead rate. We found a definite shock sensitivity, that was confined to Pol B stage 2. The shocks were applied with the "calibrated ALFA whacking tool (~1 ft length of 2x4). There was no response whatsoever to wiggling the bias cable.
Denis Urbaindurbain@naic.eduWednesday, February 13, 2013 08:27:26The C band was brought down to the lab to investigate the Pol B stage two sensitivity to induced vibrations. The Amplifier was open and inspected, a suspect solder joint on one of the resistor in the stage 2 bias circuit was found. Because of the type of solder used for the assembly of the amplifier, the solder joint could not be touched up. The resistor was replaced, and Indium solder was used to solder the new resistor that is larger than the original one. After thermaly cycling the receiver 3 times no more sensitivity to vibrations was found, and the receiver was re-installed on the antenna. The spare amplifier was sent back to Chalmers for inspection because of higher than expected noise.
Denis Urbaindurbain@naic.eduFriday, May 24, 2013 11:11:08The cold head of the C band started to show signs of contamination on the 20th of May. The cold head was turned off and the Helium lines disconnected the same day. Guarionex cleaned the coldhead that was dirty with traces of water. The receiver was cooled again on the 22nd.
Denis Urbaindurbain@naic.eduFriday, August 9, 2013 16:43:48C band was installed back on the rotary floor after the OA19 experiment ended.
Felix O. Fernandezfelix.fernandez@naic.eduThursday, February 20, 2014 10:41:28CB cold-head was turned off on 20-FEB-2014. The cold-head was emitting "knocking" sounds, its temperature was at 25K when the nominal is around 20K, and there was condensation on the Dewar, which typically indicates a loss of vacuum.
Felix O. Fernandezfelix.fernandezMonday, March 10, 2014 14:08:35Connected the vacuum pump and cold-head was replaced by ED007. The cold-head is currently off and will be turned ON once the satisfactory vacuum is achieved.
Felix O. Fernandezfelix.fernandez@naic.eduTuesday, March 11, 2014 11:37:25Vacuum reached 1 mTorr and cooling process was started.
Severe-ratchetingfelix.fernandez@naic.eduMonday, July 7, 2014 16:33:01The cold-head was ratcheting. It was decided to warm-up the receiver and service the cold-head. The displacer will be clean and baked. The cold-head cylinder will also be cleaned.
O-ring replacementfelix.fernandez@naic.eduWednesday, July 9, 2014 12:58:37An o-ring on the cold-head's supply fitting was replaced during today's work. This might have caused some He leakage.
Cold-head contaminationfelix.fernandez@naic.eduTuesday, July 29, 2014 14:59:33The CB cold-head was ratcheting and was deemed contaminated. We suspect that the contamination is due to a foreign gas since the ratcheting went off once the temperature increased some kelvins but not as much as water's freezing point (273K). We will service the cold-head tomorrow.
Re-installed Bias PCBfelix.fernandez@naic.eduTuesday, October 21, 2014 15:03:03Yesterday we took down CB's Bias PCB for examination. We found that one of the ribbon cables, specifically the one on Pol-B, was a bit loose. We re-tighten the connector and Today we replaced the Bias PCB. Hopefully this will correct the bias oscillations on Pol-B.
Pol-B bias still oscillating at nightfelix.fernandez@naic.eduTuesday, October 28, 2014 15:47:14Pol-B is still oscillating during the night. The bias oscillations on Id on Stage #2-PolB start at about 7PM and last until about 7AM. We are still baffled by this behavior.
PolB-Stage 2 Oscillationafelix.fernandez@naic.eduFriday, November 7, 2014 15:20:14PolB-Stage #2 Id current is still oscillating almost every night. Today the oscillations persisted well into the afternoon. As a side note, today has been a really rainy day. The Rx group went up to the Dome to check on CB and CBH at about 1:30PM. During our measurements we verified that the oscillations are indeed real (you could see the needle on the analog meter in the bias chasis oscillating). We tried turning the bias chasis and post-amp chasis ON and OFF trying to "kill" the oscillations but that did not work. The analog meter movements were almost rhythmic in nature so we decided to remove the cold-head connection, thinking that this might be causing the quasi-rythmic oscillations. Once we unplugged the cold-head the amplifier's Id railed out to the upper scale of the analog meter. We quickly proceeded to turn OFF the bias and post amp chasis. Once we powered them back up the oscillations were gone. We then plugged back in the cold-head and no oscillations were observed. There might be some relationship between the cold-head power cable and the oscillations were are seeing on Pol-B. We will look further into this possibility.
Pol B-Stage 2- Oscillationsfelix.fernandez@naic.eduMonday, November 10, 2014 10:49:29Today we performed a "knocking" test on CB. Basically we knocked on CB's dewar with a small piece of wood trying to get Pol B -Stage 2 to oscillate. We saw that indeed Pol B- Stage 2 did start and stopped its oscillations with very small bumps on its dewar. We can then conclude that the problem lies inside the dewar. Thus, CB must come down for repairs.
Pol-B is deadfelix.fernandez@naic.eduWednesday, November 4, 2015 11:30:11Last time CB came down we found a loose GND strap inside the dewar. After the GND strap was re-attached CB worked perfectly until last week. Suddenly, CB's PolB Stage-2 (n.b. the same stage that was having oscillations last time) went dead. We will bring down CB for repairs.
Pol-B LNA replacementfelix.fernandez@naic.eduWednesday, November 18, 2015 15:14:42Pol-B's LNA had to be replaced. The LNA would work fine at room temperature but whe it was cooled down to 20K the LNA would cease to work. At 20K the 2nd stage bias current would get "stuck" at 5.8mA and not signal output was seen at the SA. The LNA was replaced by 6H-3 18.
Cool-down started`felix.fernandez@naic.eduWednesday, October 18, 2017 08:07:47Started the CB cool-down after hurricane Maria.
Cold-head contaminationfelix.fernandez@naic.eduTuesday, September 4, 2018 15:29:31Ratcheting developed during the past weekend. Thermal cycle started this morning.
PolB High Tsys felix.fernandez@naic.eduTuesday, March 19, 2019 15:19:39Recent Pol-B's Tsys values were about 40K since 11-MAR-19 (after a DB+/-). Looking at the LNA's Id you could see that Pol-B's LNA stage 1 current was erratic during this period. After some debugging it turns out that the bias cable connecting the Bias Chassis to the dewar was loose. Today, Carlos and Wilson went up to the Dome and tighten the cable. This should fix this issue.
PolB Acting upfelix.fernandez@naic.eduTuesday, February 11, 2020 14:18:06PolB has been acting up since December 2019. Check Phil's webpage at http://www.naic.edu/~phil/calcb/polBTsys_dec19.html We did some followup and it is clear that there is something coming loose inside the Dewar. Whatever it is is coming ON and OFF at almost every cold-head cycle. We banged on the side of the dewar and signal on PolB jumped while PolA remained steady. We will thermal cycle the receiver as a stop-gap solution until we can lower the receiver.
Erratic PolBfelix.fernandez@naic.eduThursday, February 13, 2020 15:04:41The thermal cycle did NOT fix PolB's issue. We will need to bring down the receiver for repairs. PolB will remain erratic until then.
PolB LNA cryo-wire repair and internal surge protectionfelix.fernandez@naic.eduFriday, March 6, 2020 15:35:26The cryo-wire that caries the Vd voltage to stage #2 on PolB LNA had an intermittent connection at the 300k/70K terminal block. It has now been fixed. Also, we added surge protection (clamping diodes at Vd and Vg) to each LNAs' stages. Two new small PCb with the diode arrangements were added to the dewar's inside at the 300K plate. These will provide better protection to transient voltages.

Back to Top